# DecoderTracker: Decoder-Only Method for Multiple-Object Tracking

### v4

@@ Abstract
Decoder-only models, such as GPT[1], have demonstrated superior performance in many areas compared to traditional encoder-decoder structure transformer models. Over the years, end-to-end models based on the traditional transformer structure, like MOTR, have achieved remarkable performance in multi-object tracking. However, the significant computational resource consumption of these models leads to less friendly inference speeds and training times. To address these issues, this paper attempts to construct a lightweight Decoder-only model: DecoderTracker for end-to-end multi-object tracking. Specifically, drawing on some real-time detection models, we have developed an image feature extraction network which can efficiently extract features from images to replace the encoder structure. In addition to minor innovations in the network, we analyze the potential reasons for the slow training of MOTR-like models and propose an effective training strategy to mitigate the issue of prolonged training times. On the DanceTrack dataset, without any bells and whistles, DecoderTracker’s tracking performance slightly surpasses that of MOTR, with approximately twice the inference speed. Furthermore, DecoderTracker requires significantly less training time compared to MOTR.
@@ 1 Introduction
Multi-Object Tracking (MOT) is a critical task in the field of video analysis, primarily aimed at inferring and predicting the motion trajectories of objects or instances across a sequence of continuous images. The success of transformer[2] models in various domains in recent years is well-documented, and recently, transformer-based end-to-end MOT models like MOTR[3] have been introduced in the MOT domain. These models offer advantages over the widely used tracking-by-detection methods[4, 5, 6, 7], such as eliminating the need for manual design or selection of features. Additionally, some enhanced methods[8, 9, 10] of MOTR have demonstrated the potential of this model class in terms of tracking accuracy.
However, the inference speed of these models is often suboptimal, and their RNN[11] like linear inference structure makes them less GPU friendly during training.
@@
At the same time, we discovered that the original training process for MOTR is one of strong supervision, which led us to question whether using relatively weaker supervision and allowing the model to engage in less stringent learning might also be effective. Is there a possibility that this process could yield unexpected benefits? Based on this idea, we proposed the Tracking Box Selection Process (TBSP).
Although the methods we propose have not yet enabled the inference speed of MOTR-like models to achieve real-time performance, we believe that this paper could provide valuable insights for future researchers.
@@
    A second key innovation introduces a unique training strategy characterized by a three-stage process. This strategy accelerates model convergence, reducing overall training time and enhancing training process efficiency.
@@
    The third contribution introduces TBSP, a strategy that can serve as a preliminary weakly supervised training approach for end-to-end MOT model training.
@@ 2 Related work
Figure 1: MOTR series’ Streamline the process of updating detect queries and track queries under typical MOT scenarios. The track query set, initially empty, is dynamically adjusted in length. Detect queries play a role in identifying new objects.
@@ 2.1 MOTR Series
Unlike tracking-by-detection models exemplified by SORT[4, 14, 5], and previous transformer-based MOT models[15, 16], MOTR[3] is the first model to achieve end-to-end multi-object tracking. It extends the object query from DETR[17] into a track query, thereby enabling tracking capabilities. It uses a learnable positional embedding and multi-scale deformable attention to predict object positions and classes in frames, eliminating the need for explicit data association or post-processing. Key innovations include the Track Query Aggregation with Learned Attention (TALA), which uses a query memory bank and multi-head attention for modeling long-term temporal relationships, and the introduction of Collective Average Loss (CAL) and Temporal Aggregation Network (TAN) to resolve target conflicts and improve temporal information integration. These features establish MOTR as a significant advancement in multi-object tracking technology.
@@
Similarly, MeMOT[10] innovates by incorporating track queries that model and predict target trajectories iteratively across video frames. This approach utilizes a tracklet-aware label assignment method to ensure a one-to-one correspondence between track queries and actual target trajectories, addressing challenges with newly introduced object queries.
Moreover, MOTRv3[9] and CO-MOT[20], as detailed in their respective papers, focus on rectifying the imbalance in label assignment between detection and tracking queries within the MOTR framework. MOTRv3, adopts a release-fetch supervision strategy to facilitate balanced label assignments, enhancing overall MOT performance. In contrast, CO-MOT, employs a coopetition label assignment strategy that not only increases the number of positive samples for detection queries but also introduces "shadows" to enhance the diversity and quantity of queries, further refining the MOT capabilities.
@@ 2.2 Real Time Detection Method
When discussing real-time detection models, most people initially think of YOLO[21], which transformed the problem of object detection into a regression model, achieving efficient object detection. The model divides an image into grids, with each grid responsible for predicting the bounding boxes and class information of objects within its area, treating the task as a regression problem of coordinates and dimensions. The most advanced versions in the YOLO series currently are YOLOv8.2[22] and YOLOv9[23].
Recently, the real-time detection model based on DETR, known as RT-DETR[24], has also demonstrated its extraordinary performance. RT-DETR achieves optimized inference speeds by designing an efficient encoder and simplifying the model structure.
Although we acknowledge the efficiency of RT-DETR’s hybrid encoder, we believe that the encoder structure might be unnecessary. Therefore, inspired by the YOLO series[25, 22, 26], we designed a feature extraction network.
3 DecoderTracker
Figure 2: The overall structure of DecoderTracker. By merging the detect queries ($q_{d}$) and track queries ($q_{tr}$), the decoder (Dec) processes them to generate hidden states. When the model is trained using the TALA strategy, the TBSP can be skipped.
Figure 3: The structure of IMFENet. It bears certain similarities to the structure of YOLOv8’s backbone and neck, yet there are distinct differences. For detailed comparisons, please refer to the appendix.
3.1 DecoderTracker Architecture
DecoderTracker’s overall architecture is illustrated in the provided Fig.2. In this setup, a video sequence is input to Image Feature Extraction Network(IMFENet) to extract features. For the first frame, where there is no tracking information available, a fixed-length learnable detect query (referred to as $q_{d}$ in the figure) is input to the decoder.
@@ 3.1 DecoderTracker Architecture
The structure of IMFENet is primarily inspired by YOLOv8, which replaces the backbone and encoder structures of the original MOTR for image feature extraction. The reason for not adopting the latest YOLOv9 is due to its slower training speed compared to YOLOv8. Through IMFENet, we encode an image of dimensions $H\times W\times 3$ into three tensors of sizes $\frac{H}{8}\times\frac{W}{8}\times 256$,$\frac{H}{16}\times\frac{W}{16}\times 256$, $\frac{H}{32}\times\frac{W}{32}\times 256$, which are then fed into the decoder.
3.2 Tracking Box Selection Process
Figure 4: The specific process of TBSP is represented by a formula. The specific meaning of the formulas in the figure refers to Section 3.2
@@ 3.2 Tracking Box Selection Process
Consequently, we designed a simple filtering mechanism, TBSP, which enables the detection network to possess some tracking capabilities without additional training. However, given the network’s inherently weak tracking ability and the lack of robustness in TBSP, the tracking performance at this stage is quite limited. Nevertheless, we also discovered that using TBSP as a preprocessing step for TALA, and training the network in this fashion, can lead to some unexpected benefits.
Compared to TALA, TBSP can be considered a form of self-supervised or weak supervision during training. This is because TBSP does not forcibly assign a specific query to correspond to an object; instead, it allows the model to optimize itself. For a discussion on why TBSP is effective, please refer to the appendix.
We know that both track query and detect query contain information about the object’s bounding box(BBox). Let’s denote the BBox in $\hat{\omega}_{tr}^{t-1}$ as $[BBox_{1}^{tr}$, $BBox_{2}^{tr},...,BBox_{n}^{tr}]$, where $n$ is the number of track queries at time $t-1$. Each query in $\hat{\omega}_{tr}^{t-1}$ is denoted as $[\omega_{1}^{det},\omega_{2}^{det},...,\omega_{m}^{det}]$, and its corresponding BBox is denoted as $[BBox_{1}^{det},BBox_{2}^{det},...,BBox_{m}^{det}]$.
@@
The feasibility of TBSP is illustrated in the Fig. 4. Incidentally, although TBSP makes DecoderTracker appear not to be a truly end-to-end network, after thorough TALA training, this step can be omitted. Simultaneously, to better illustrate the performance of our model compared to MOTR, we have not adopted the labeling strategy used in CO-MOT[8] and MOTRv3[9].
3.3 Loss and Training Methods
Due to MOTR’s similarity to RNN in its inference approach, its batch size is limited to 1. This often results in suboptimal utilization of GPU resources during training. Additionally, MOTR learns both image and motion features simultaneously. To expedite model training, this paper adopts a three-stage training strategy. In the first stage, emphasis is placed on training the network to extract target image features efficiently. The second stage involves training the model on weak tracking aspects using TBSP as a pretraining step for the third stage. Lastly, the third stage employs the TALA strategy to comprehensively enhance the model’s tracking performance.
@@ 3.3 Loss and Training Methods
$\mathcal{L}_{o}\left(\left.\widehat{Y}\right|_{\omega},Y\right)=\frac{\mathcal{L}\left(\left.\widehat{Y}_{\det}^{t}\right|_{\omega_{\det}^{t}},Y_{\det}^{t}\right)}{V_{i}}$ (2)
$\mathcal{L}$ is the loss of single frame. Notably, during this stage, we employed the IoU-aware Query Selection from RT-DETR, so the loss function utilized is from RT-DETR. Unlike the original DETR, this stage introduces a loss termed $\mathcal{L}_{enc}$. This loss involves treating the initialized queries as boxes for the purpose of loss computation. The specifics are outlined below:
$\mathcal{L}_{enc}(\hat{y},y)&=\mathcal{L}_{\text{box }}(\hat{b},b)+\mathcal{L}_{cls}(\hat{c},\hat{b},y,b)\\ &=\mathcal{L}_{\text{box }}(\hat{b},b)+\mathcal{L}_{cls}(\hat{c},c,IoU)\end{aligned}$ (3)
where $\hat{y}$ and $y$ denote prediction and ground truth, $\hat{y}={\hat{c},\hat{b}}$ and $y={c,b}$, $c$ and $b$ represent categories and bounding boxes, respectively. Finally, the overall loss is as follows:
@@
where $\hat{\omega}_{tr}^{t-1}$ represents the multi-frame predictions,$\hat{Y}=\left\{\hat{Y}_{i}\right\}_{i=1}^{N}$ represents the loss calculated over the entire video sequence, using ground truth $Y=\left\{Y_{i}\right\}_{i=1}^{N}$ and matched results $\omega=\left\{\omega_{i}\right\}_{i=1}^{N}$ $\mathcal{L}_{cls}$ denotes the focal loss[27], $\mathcal{L}_{cls}$ represents the L1 loss and $\mathcal{L}_{giou}$ represents the giou loss[28], $\lambda_{cls},\lambda_{l1},\lambda_{giou},\lambda_{enc}$ are corresponding weight coefficients.
In this stage, the network is capable of initially learning the appearance features of objects to be tracked. During this phase, the batch size can be set larger, fully utilizing the computational power of GPUs to reduce training time. Additionally, various data augmentation techniques [29, 30, 31] can be employed to expedite the model’s extraction of appearance features.
@@
$\mathcal{L}_{o}\left(\left.\widehat{Y}\right|_{\omega},Y\right)=\frac{\sum_{n=1}^{N}{\left(\mathcal{L}\left(\left.\widehat{Y}_{tr}^{t}\right|_{\omega_{tr}^{t}},Y_{tr}^{t}\right)+\mathcal{L}\left(\left.\widehat{Y}_{\det}^{t}\right|_{\omega_{\det}^{t}},Y_{\det}^{t}\right)\right)}}{\sum_{n=1}^{N}{\left(V_{t}\right)}},N=5$ (5)
@@
where $V_{i}=V_{tr}^{i}+V_{det}^{i}$ denotes the total number of ground-truths objects at frame $i$. $V_{tr}^{i}$ and $V_{det}^{i}$ are the numbers of tracked objects and newborn objects at frame i, respectively.
@@
At this stage, the computation of $\mathcal{L}_{enc}$ also differs from that in the first phase, where the newly generated detection queries are calculated only in conjunction with newly appeared objects. Moreover, in $\mathcal{L}\left(\left.\widehat{Y_{tr}^{t}}\right|_{\omega_{tr}^{t}},Y_{tr}^{t}\right)$, $\mathcal{L}_{enc}$ is absent, and the weight of $\lambda_{enc}$ is reduced compared to it in second stage.

### v5

# DecoderTracker: Decoder-Only End-To-End method for Multiple-Object Tracking

@@ Abstract
Decoder-only methods, such as GPT, have demonstrated superior performance in many areas compared to traditional encoder-decoder structure transformer methods. Over the years, end-to-end methods based on the traditional transformer structure, like MOTR, have achieved remarkable performance in multi-object tracking. However,The substantial computational resource consumption of these methods, coupled with the optimization challenges posed by dynamic data, results in less favorable inference speeds and training times. To address the aforementioned issues, this paper optimized the network architecture and proposed an effective training strategy to mitigate the problem of prolonged training times, thereby developing DecoderTrack, a novel end-to-end tracking method. Subsequently, to tackle the optimization challenges arising from dynamic data, this paper introduced DecoderTrack+ by incorporating a Fixed-Size Query Memory and refining certain attention layers. Our methods, without any bells and whistles, outperforms MOTR on multiple benchmarks, with inference speeds 2.06 and 3.03 times faster than MOTR, respectively.
keywords:
Multi-object tracking , Decoder , End-to-end
journal: Pattern Recognition11footnotetext: This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
@@ 1 Introduction
Multi-Object Tracking (MOT) is a critical task in the field of video analysis, primarily aimed at inferring and predicting the motion trajectories of objects or instances across a sequence of continuous images. The success of transformer [1] methods in various domains in recent years is well-documented, and recently, transformer-based end-to-end MOT methods like MOTR [2] have been introduced in the MOT domain. These methods offer advantages over the widely used tracking-by-detection methods [3, 4, 5, 6], such as eliminating the need for manual design or selection of features. Additionally, some enhanced methods [7, 8, 9] of MOTR have demonstrated the potential of this method class in terms of tracking accuracy.
@@
Figure 1: MOTR series’ Streamline the process of updating detect queries and track queries under typical MOT scenarios. The number of detect queries is typically fixed, while the number of track queries fluctuates throughout the tracking process. This discrepancy results in dynamic data during inference, which introduces optimization challenges that limit the actual runtime speed of the method. To address this, we propose the Fixed-Size Query Memory (FSQM), which pads the tracking query sequence with empty queries to maintain a consistent number of queries during inference. By eliminating dynamic data, we have developed DecoderTracker+, which has significantly improved inference speed compared to MOTR and DecoderTracker.
At the same time, we discovered that the original training process for MOTR is one of strong supervision, which led us to question whether using relatively weaker supervision and allowing the method to engage in less stringent learning might also be effective. Is there a possibility that this process could yield unexpected benefits? It is important to note that weak supervision, or self-supervised learning, has achieved significant success in many computer vision tasks [12, 13]. Based on this idea, we proposed the Tracking Box Selection Process (TBSP) which is mainly utilized during the training. Thus, DecoderTracker has been constructed by us.
After completing the aforementioned architectural optimizations mentioned above, we observed that while the training duration was significantly optimized, the inference speed did not improve as expected. This phenomenon prompted us to delve deeper into the issue. In theory, streamlining the network architecture should more notably enhance inference speed rather than training efficiency. To investigate this anomaly, we systematically analyzed the time consumption distribution across various modules. The results revealed that the decoder module accounted for the majority of the time, consuming nearly two-thirds of the model’s total inference time, which is evidently unreasonable. Through further cross-model comparative analysis, we identified that this issue also exists in MOTR but is absent in its base model, Deformable DETR. Controlled variable experiments indicated that fixing the number of queries significantly reduces decoder latency. Considering factors such as GPU memory usage, we ultimately pinpointed the root cause to be the dynamic query mechanism. This mechanism introduces unnecessary overhead during inference, such as inefficient GPU memory allocation in modern deep learning frameworks ,e.g., PyTorch/TensorFlow, and may also hinder compilation optimizations like operator fusion due to dynamic computation graphs. To address this issue, we proposed the Fixed-Size Query Memory (FSQM) for handling fixed-query tracking. By integrating FSQM and optimizing the attention layers in the original decoder and TAN modules, we introduced DecoderTrack+. Experimental results demonstrated that, although DecoderTrack+ theoretically increases computational load, its inference speed is significantly faster than the original version (from 19.6 fps to 28.8 fps) due to static data processing, while maintaining comparable tracking performance.
@@
    The second key innovation involves the introduction of a unique training strategy characterized by a three-stage process. At the same time, we incorporated TBSP (Temporal Balanced Sampling and Propagation), a weakly supervised training strategy designed for preliminary training of end-to-end MOT methods. These strategies accelerate method convergence, reduce overall training time, and enhance the efficiency of the training process.
@@
    Finally, by leveraging FSQM and optimizing certain self-attention layers within the network, we developed DecoderTrack+, which addresses the additional latency caused by dynamic data. This provides a fully viable technical pathway for the engineering deployment of MOTR-like end-to-end MOT models.
@@ 2.1 MOTR Series
Unlike tracking-by-detection methods exemplified by SORT [3, 14, 4], and previous transformer-based MOT methods [15, 16], MOTR [2] is the first method to achieve end-to-end multi-object tracking. It extends the object query from DETR [17] into a track query, thereby enabling tracking capabilities. It uses a learnable positional embedding and multi-scale deformable attention to predict object positions and classes in frames, eliminating the need for explicit data association or post-processing. Key innovations include the Track Query Aggregation with Learned Attention (TALA), which uses a query memory bank and multi-head attention for methoding long-term temporal relationships, and the introduction of Collective Average Loss (CAL) and Temporal Aggregation Network (TAN) to resolve target conflicts and improve temporal information integration. These features establish MOTR as a significant advancement in multi-object tracking technology.
@@
Similarly, MeMOT [9] innovates by incorporating track queries that method and predict target trajectories iteratively across video frames. This approach utilizes a tracklet-aware label assignment method to ensure a one-to-one correspondence between track queries and actual target trajectories, addressing challenges with newly introduced object queries.
MOTRv3 [8] and CO-MOT [20], as detailed in their respective papers, focus on rectifying the imbalance in label assignment between detection and tracking queries within the MOTR framework. MOTRv3, adopts a release-fetch supervision strategy to facilitate balanced label assignments, enhancing overall MOT performance. In contrast, CO-MOT, employs a coopetition label assignment strategy that not only increases the number of positive samples for detection queries but also introduces ”shadows” to enhance the diversity and quantity of queries, further refining the MOT capabilities.
2.2 Tracking by Detection
Tracking by Detection are among the most classical and widely used types of MOT. These methods typically employ a detector to first identify objects and then use various other methods for tracking.
In past years, the main methods within the Tracking by Detection category were SORT [3] and its derivative methods. The primary approach of these methods is to use a Kalman Filter to predict the positions of targets in the next frame and then perform data association with the detection results provided by the detector. Many subsequent improvements to these methods [5, 14, 4, 21] have introduced re-identification (reid) networks, incorporating object appearance features as an additional factor for association. There are also a series of methods like OC-SORT [22] that purely improve the Kalman Filter-based tracking method using kinematic methods.
In recent years, methods such as SUSHI [23], Glan [24] and PuTR [25] have emerged, which employ networks for tracking post-detection. The former two utilize Graph Neural Networks (GNN), while the last is based on transformer.
@@ 2.3 Real Time Detection Method
Figure 2: The overall structure of DecoderTracker. By merging the detect queries ($q_{d}$) and track queries ($q_{tr}$), the decoder (Dec) processes them to generate hidden states. To put it simply, $q_{tr}$ represents the objects in the video sequence that need to be tracked continuously, while $q_{d}$ represents the new objects appearing in the video frame. When the method is trained using the TALA strategy, the TBSP can be skipped. The process of DecoderTracker+ is consistent with this scenario, but the main difference lies in the updating of $q_{tr}$ and whether it is of fixed length. The FSQM of DecoderTracker+ pads $q_{tr}$ with all-zero queries. This approach effectively addresses optimization issues caused by dynamic data.
When discussing real-time detection methods, most people initially think of YOLO [26], which transformed the problem of object detection into a regression method, achieving efficient object detection. The method divides an image into grids, with each grid responsible for predicting the bounding boxes and class information of objects within its area, treating the task as a regression problem of coordinates and dimensions. The most advanced versions in the YOLO series currently are YOLOv8.2 [27] and YOLOv9 [28].
Recently, the real-time detection method based on DETR, known as RT-DETR [29], has also demonstrated its extraordinary performance. RT-DETR achieves optimized inference speeds by designing an efficient encoder and simplifying the method structure.
Although we acknowledge the efficiency of RT-DETR’s hybrid encoder, we believe that the encoder structure might be unnecessary. Therefore, inspired by the YOLO series [28, 27], we designed a feature extraction network.
3 Method
3.1 DecoderTracker
Figure 3: The structure of FENet. It bears certain similarities to the structure of YOLOv8’s backbone and neck, yet there are distinct differences. Specific module details can be found in Figure 4. The extracted features will be fed into the Decoder and introduced to the cross-attention module to interact with the queries. Similar to the Backbone and Encoder in the MOTR series, FENet focuses solely on extracting features from the current video frame, while the implicit spatio-temporal feature extraction is managed by the Decoder and the TAN.
3.1.1 DecoderTracker Architecture
DecoderTracker’s overall architecture is illustrated in the provided Figure 2. In this setup, a video sequence is input to image Feature Extraction Network(FENet) to extract features. For the first frame, where there is no tracking information available, a fixed-length learnable detect query (referred to as $q_{d}$ in the figure) is input to the decoder.
@@ 3.1.1 DecoderTracker Architecture
3.1.2 Design of FENet
The structure of FENet is primarily inspired by YOLOv8, which replaces the backbone and encoder structures of the original MOTR for image feature extraction. The reason for not adopting the latest YOLOv9 or YOLOv10 is due to its slower training speed compared to YOLOv8. Through FENet, we encode an image of dimensions $H\times W\times 3$ into three tensors of sizes $\frac{H}{8}\times\frac{W}{8}\times 256$,$\frac{H}{16}\times\frac{W}{16}\times 256$, $\frac{H}{32}\times\frac{W}{32}\times 256$, which are then fed into the decoder.
The design of this FENet is intended to enhance the method’s efficiency and increase its inference speed. Thus, our design primarily references the network structure of YOLOv8. The backbone structure of our method is almost identical to that of YOLOv8. Our main modification involves changes to the neck structure in YOLOv8, to ensure that its output aligns with the input requirements of the decoder.
Figure 4: Specific module details, modules not specified in the diagram are network structures or APIs that come with PyTorch. These modules are primarily constructed based on CNN, capable of efficiently extracting image features. The employment of residual connections ensures that issues like gradient explosion and vanishing gradients are largely avoided.
3.1.3 Tracking Box Selection Process
Algorithm 1 TBSP flow
0:  $\hat{\omega}_{tr}^{t-1}$: Track queries assignment, $\hat{\omega}_{det}^{t-1}$: Detect queries assignment, $\delta_{iou}$: IOU threshold
1:  $[BBox_{1}^{tr},BBox_{2}^{tr},\dots,BBox_{n}^{tr}]\leftarrow$ BBoxes from $\hat{\omega}_{tr}^{t-1}$
2:  $[BBox_{1}^{det},BBox_{2}^{det},\dots,BBox_{m}^{det}]\leftarrow$ BBoxes from $\hat{\omega}_{det}^{t-1}$
3:  for  $BBox_{j}^{det}$ in $[BBox_{1}^{det},\dots,BBox_{m}^{det}]$ do
4:     for  $BBox_{i}^{tr}$ in $[BBox_{1}^{tr},\dots,BBox_{n}^{tr}]$ do
5:        Compute $IOU(BBox_{j}^{det},BBox_{i}^{tr})$
6:        if $IOU>\delta_{iou}$ then
7:           Discard $BBox_{j}^{det}$
8:           break
9:        end if
10:     end for
11:  end for
12:  $\omega_{\det_{f}}^{t-1}\leftarrow$ Remaining detect queries assignment
@@ 3.1.3 Tracking Box Selection Process
Consequently, we designed a simple filtering mechanism, TBSP, which enables the detection network to possess some tracking capabilities without additional training. However, given the network’s inherently weak tracking ability and the lack of robustness in TBSP, the tracking performance at this stage is quite limited. Nevertheless, we also discovered that using TBSP as a preprocessing step for TALA, and training the network in this fashion, can lead to some unexpected benefits. Compared to TALA, TBSP can be considered a form of self-supervised or weak supervision during training. This is because TBSP does not forcibly assign a specific query to correspond to an object; instead, it allows the method to optimize itself.
Formally, we denote the predictions of track queries as $\widehat{Y}_{tr}$ and predictions of detect as $\widehat{Y}_{\text{det }}.Y_{\text{new }}$ is the ground-truths of newborn objects. The label assignment results for track queries and detect queries can be written as $\omega_{tr}$ and $\omega_{\text{det}}$. For frame $i$, label assignment for detect queries is obtained from bipartite matching among detect queries and newborn objects. We know that both track query and detect query contain information about the object’s bounding box (BBox). Let’s denote the BBox in $\hat{\omega}_{tr}^{t-1}$ as $[BBox_{1}^{tr}$, $BBox_{2}^{tr},...,BBox_{n}^{tr}]$, where $n$ is the number of track queries at time $t-1$. Each query assignment in $\hat{\omega}_{det}^{t-1}$ is denoted as $[\omega_{1}^{det},\omega_{2}^{det},...,\omega_{m}^{det}]$, and its corresponding BBox is denoted as $[BBox_{1}^{det},BBox_{2}^{det},...,BBox_{m}^{det}]$.
@@
The feasibility of TBSP is illustrated in the Algorithm. 1. Incidentally, although TBSP makes DecoderTracker appear not to be a truly end-to-end network, after thorough TALA training, this step can be omitted. Simultaneously, to better illustrate the performance of our method compared to MOTR, we have not adopted the labeling strategy used in CO-MOT [7] and MOTRv3 [8].
Figure 5: Refinement process for DecoderTracker on partially trained track queries. When DecoderTracker undergoes comprehensive training with the TALA strategy, TBSP will no longer be utilized. At that point, the track query generation method of DecoderTracker aligns with MOTR, as illustrated in the Figure. 1. The process in DecoderTracker+ is similar, but there are some differences in the removal and addition of track queries. The detailed operations are discussed in the subsequent sections.
3.1.4 Query and Query Positions
MOTR was initially built upon Deformable DETR [11]. The original DETR was considered slow to converge due to its Hungarian matching algorithm. Although Deformable DETR significantly improved the convergence speed of DETR-like methods through architectural enhancements, subsequent works such as DN-DETR [30] and DAB-DETR [31] further corroborated the assertion that the Hungarian matching algorithm was indeed problematic. In DETR and Deformable DETR, the queries are typically initialized as a set of non-interpretable parameters, which is evidently suboptimal for training. Drawing inspiration from RT-DETR’s query generation approach, we directly generate queries from the input decoder features f. This process can be specifically represented as follows:
$q_{d}&=\text{LayerNorm}(\text{Linear}(f))\\ \end{aligned}$ (2)
One more thing, to expedite method convergence, we implement a Query Positions generation strategy akin to CO-MOT. Drawing inspiration from DAB-DETR [31], we employ a fixed transformation to convert reference points into Query Positions. While experimentation with the RT-DETR approach involving a Multilayer Perceptron (MLP) to transform reference points shows no significant performance discrepancy, it is important to note that the use of MLP substantially increases training time.
3.1.5 Loss and Training Methods
To expedite method training, this paper adopts a three-stage training strategy. In the first stage, emphasis is placed on training the network to extract target image features efficiently. The second stage involves training the method on weak tracking aspects using TBSP as a pretraining step for the third stage. Lastly, the third stage employs the TALA strategy to comprehensively enhance the method’s tracking performance.
@@ 3.1.5 Loss and Training Methods
where $\hat{Y}=\left\{\hat{Y}_{i}\right\}_{i=1}^{N}$ represents the frame predictions. The loss is calculated by ground truth $Y=\left\{Y_{i}\right\}_{i=1}^{N}$ and matched results $\omega=\left\{\omega_{i}\right\}_{i=1}^{N}$. $\mathcal{L}_{cls}$ denotes the focal loss[32], $\mathcal{L}_{cls}$ represents the L1 loss and $\mathcal{L}_{giou}$ represents the giou loss[33]. $\mathcal{L}_{enc}$ is the anchor loss of DAB-DETR [31] and $\lambda_{cls},\lambda_{l1},\lambda_{giou},\lambda_{enc}$ are corresponding weight coefficients, which are set to 1,2,5,5 in our method, just like DAB-DETR.
In this stage, the network is capable of initially learning the appearance features of objects to be tracked. During this phase, the batch size can be set larger, fully utilizing the computational power of GPUs to reduce training time. Additionally, various data augmentation techniques [34, 35, 36] can be employed to expedite the method’s extraction of appearance features.
@@
[…]
@@
where $V_{i}=V_{tr}^{i}+V_{det}^{i}$ denotes the total number of ground-truths objects at frame $i$. $V_{tr}^{i}$ and $V_{det}^{i}$ are the numbers of tracked objects and newborn objects at frame i, respectively. $\mathcal{L}$ is the loss of single frame, which is similar to Eq.3.
@@
At this stage, the computation of $\mathcal{L}_{enc}$ also differs from that in the first phase, where the newly generated detection queries are calculated only in conjunction with newly appeared objects. Moreover, in $\mathcal{L}\left(\left.\widehat{Y_{tr}^{t}}\right|_{\omega_{tr}^{t}},Y_{tr}^{t}\right)$, $\mathcal{L}_{enc}$ is absent, and the weight of $\lambda_{enc}$ is reduced compared to it in second stage. Thanks to our various strategies and improvements to the network, we achieved huge improvements in training time and inference speed compared to MOTR.